### v2

# Pychop: Emulating Low-Precision Arithmetic in Numerical Methods and Neural Networks

Abstract.
Motivated by the growing demand for low-precision arithmetic in computational science, we exploit lower-precision emulation in Python—widely regarded as the dominant programming language for numerical analysis and machine learning. Low-precision training has revolutionized deep learning by enabling more efficient computation and reduced memory and energy consumption while maintaining model fidelity. To better enable numerical experimentation with and exploration of low precision computation, we developed the Pychop library, which supports customizable floating-point formats and a comprehensive set of rounding modes in Python, allowing users to benefit from fast, low-precision emulation in numerous applications. Pychop also introduces interfaces for both PyTorch and JAX, enabling efficient low-precision emulation on GPUs for neural network training and inference with unparalleled flexibility.
In this paper, we offer a comprehensive exposition of the design, implementation, validation, and practical application of Pychop, establishing it as a foundational tool for advancing efficient mixed-precision algorithms. Furthermore, we present empirical results on low-precision emulation for image classification and object detection using published datasets, illustrating the sensitivity of the use of low precision and offering valuable insights into its impact. Pychop enables in-depth investigations into the effects of numerical precision, facilitates the development of novel hardware accelerators, and integrates seamlessly into existing deep learning workflows. Software and experimental code are publicly available at https://github.com/inEXASCALE/pychop.
Mixed Precision Simulation, Python, Neural Networks, Quantization, Numerical Methods, Deep Learning
The first author was funded by the Charles University Research Centre program (No. UNCE/24/SCI/005). The first and second authors were funded by the European Union (ERC, inEXASCALE, 101075632). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
ccs: Mathematics of computing Mathematical software performanceccs: Computing methodologies Modeling and simulationccs: Software and its engineering
1. Introduction
The advent of hardware architectures that natively support low-precision arithmetic has catalyzed a resurgence of interest in mixed-precision algorithms, particularly within the fields of numerical linear algebra and deep neural network (DNN) training. These algorithms, which strategically integrate low-precision and high-precision computations, have emerged as a focus of research due to their capacity to optimize algorithmic performance across various areas—most notably energy efficiency, computational speedup, and resource utilization—while maintaining acceptable levels of numerical accuracy and stability (Higham and Mary, 2022). This paradigm exploits the inherent trade-offs between computational cost and precision, offering a compelling framework for addressing the escalating demands of large-scale scientific simulations and machine learning applications. The ability to tailor precision to specific computational tasks not only reduces power consumption but also reduces processing times, making mixed-precision techniques indispensable in the era of exascale computing and resource-constrained environments.
The advance of mixed-precision training has been a milestone for deep learning efficiency. Micikevicius et al. (2018) pioneered its practical application, demonstrating that fp16 (half-precision, 16 bits total with 5 exponent and 10 significand bits) could handle weights and activations while fp32 accumulations preserved gradient fidelity, yielding speedups of 23×2-3$\times$2 - 3 × on NVIDIA Volta GPUs. This approach has been standardized in PyTorch (Paszke et al., 2019) via Automatic Mixed Precision (AMP) [^1] https://pytorch.org/docs/stable/amp.html, which automates precision switching and gradient scaling, reducing memory usage by approximately 50% for models like ResNet-50 (He et al., 2016). However, AMP’s dependence on hardware-supported fp16 limits its adaptability to other formats, a constraint our emulator eliminates by supporting arbitrary precision configurations.
Central to this research domain is the advancement of quantization methodologies, encompassing binary, ternary, and 4- to 8-bit schemes, which facilitate substantial compression of model parameters and intermediate representations with minimal degradation in performance (Jacob et al., 2018). Such techniques are pivotal for reducing memory footprints and computational overhead, rendering them especially valuable for deploying DNNs on edge devices and other hardware with limited resources. To preserve the integrity of the compressed information, it is essential to devise quantization strategies that explicitly maintain similarity between the original and quantized representations, a challenge that necessitates sophisticated similarity-preserving algorithms (Zhu et al., 2021). Nevertheless, the deployment of large language models (LLMs) at scale introduces additional complexities, as achieving high availability and scalability remains a significant hurdle. Overcoming these obstacles requires innovative solutions in lexical search, data re-ranking, and retrieval-augmented generation, all optimized to function efficiently at reduced computational costs (Lewis et al., 2020). The successful integration of these techniques not only enhances the efficacy of LLMs but also broadens their applicability, enabling real-time inference and supporting data-intensive research endeavors across diverse domains. As the field progresses, the synergy between mixed-precision computing and quantization promises to redefine the boundaries of computational efficiency and model performance, paving the way for more sustainable and accessible artificial intelligence systems.
Table 1. Key parameters of six floating point formats; $u$ denotes the unit roundoff corresponding to the precision, $x_{\min}$ denotes the smallest positive normalized floating-point number, $x_{\max}$ denotes the largest floating-point number, $t$ denotes the number of binary digits in the significand (including the implicit leading bit), $e_{\min}$ denotes exponent of $x_{\min}$, $e_{\max}$ denotes exponent of $x_{\max}$, with added columns for exponent bits and significand bits (excluding implicit bit).
$\quad u$ $\quad x_{\min}$ $\quad x_{\max}$ $t$ $e_{\min}$ $e_{\max}$ exp. bits sig. bits
NVIDIA quarter precision (q43) $6.25\times 10^{-2}$ $1.5625\times 10^{-2}$ $2.40\times 10^{2}$ 4 -6 7 4 3
NVIDIA quarter precision (q52) $1.25\times 10^{-1}$ $6.10\times 10^{-5}$ $5.73\times 10^{4}$ 3 -14 15 5 2
bfloat16 (bf16) $3.91\times 10^{-3}$ $1.18\times 10^{-38}$ $3.39\times 10^{38}$ 8 -126 127 8 7
half precision (fp16) $4.88\times 10^{-4}$ $6.10\times 10^{-5}$ $6.55\times 10^{4}$ 11 -14 15 5 10
TensorFloat-32 (tf32) $9.77\times 10^{-4}$ $1.18\times 10^{-38}$ $1.70\times 10^{38}$ 11 -126 127 8 10
single (fp32) $5.96\times 10^{-8}$ $1.18\times 10^{-38}$ $3.40\times 10^{38}$ 24 -126 127 8 23
double (fp64) $1.11\times 10^{-16}$ $2.23\times 10^{-308}$ $1.80\times 10^{308}$ 53 -1022 1023 11 52
Our work addresses this limitation with a mixed-precision emulation software module for Python, offering a highly configurable platform to simulate arbitrary low precision floating-point formats and a diverse array of rounding modes. This emulator transcends the constraints of fixed hardware by allowing users to define custom precision configurations—specifying exponent and significand bits—and to select from rounding modes: round to nearest, round up / down, round toward zero, round toward odd, and two stochastic variants (proportional and uniform probability). Its advantages are substantial and multifaceted:
  • Unmatched Flexibility: Emulates standard formats (see Table 1) and custom designs, enabling researchers to prototype hypothetical hardware or explore theoretical precision limits without physical constraints.
  • Precision Granularity: Provides precise control over numerical representation, critical for dissecting the impact of quantization on gradient updates, weight distributions, and activation ranges in neural networks.
  • Seamless Integration: Embeds directly into PyTorch layers, allowing practitioners to experiment with mixed precision in production-grade training pipelines with minimal overhead.
  • Rounding Mode Exploration: Supports a comprehensive set of rounding strategies, facilitating detailed studies of their effects on numerical stability, convergence rates, and final model performance—an underrepresented area in hardware-based implementations.
This paper is organized as follows: Section 2 reviews prior work on precision emulation software, identifying gaps that our solution addresses; Section 3 describes the implementation and usage in detail; and Section 4 presents simulated experiments that collectively demonstrate the emulator’s performance and its value in advancing mixed-precision emulation for deep learning. Section 5 concludes the paper and outlines future work.
2. Related Work
Table 2. Software for Simulating Low-precision Arithmetic
Package name Primary language Storage format Target format Rounding modes FPQ IQ NN STE
$\mathbf{p}$ $\mathbf{e}$ $\mathbf{s}$ built-in
RNE
RNZ
RNA
RZ
RUD
RO
SR
GNU MPFR (Fousse et al., 2007) C custom A A O $\checkmark$ ! $\checkmark$ $\checkmark$
SIPE (Lefèvre, 2013) C multiple R S Y $\checkmark$ $\checkmark$
rpe (Dawson and Düben, 2017) Fortran fp64 R B B fp16 $\checkmark$
FloatX (Flegar et al., 2019) C++ fp32/fp64 R S Y $\checkmark$
FlexFloat (Tagliavini et al., 2020) C++ fp32/fp64 R S Y $\checkmark$
INTLAB (Rump, 1999) MATLAB fp64 R S Y $\checkmark$ $\checkmark$ $\checkmark$
chop (Higham and Pranesh, 2019) MATLAB fp32/fp64 R S F fp16/bf16 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
QPyTorch (Zhang et al., 2019) Python fp32 R S N $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CPFloat (Fasi and Mikaitis, 2023) C fp32/fp64 R S F fp16/bf16/tf32 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Pychop Python / MATLAB fp32/fp64 R S F fp16/bf16 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
The columns are categorized as follows: (i) Package name, Primary language, Storage format: The name of the package, its primary programming language, and the supported storage formats. (ii) Target format parameters: $p$: Number of bits of precision in the significand—arbitrary (A) or restricted to the storage format’s significand (R); $e$: Exponent range—arbitrary (A), a sub-range of the storage format (S), or a sub-range only for built-in types (B); $s$: Support for subnormal numbers—supported (Y), not supported (N), supported only for built-in types (B), supported by default but can be disabled (F), or not supported by default but can be enabled (O); built-in: Floating-point formats natively built into the system (e.g., fp16, bf16, tf32). (iii) Rounding modes: Supported modes include round-to-nearest with ties-to-even (RNE), ties-to-zero (RNZ), ties-to-away (RNA), round-toward-zero (RZ), round up/down (RUD), round-to-odd (RO), and stochastic rounding variants (SR). A $\checkmark$ indicates full support, while ! denotes modes that can be simulated at a higher computational cost. (iv) FPQ: support fixed-point quantization. (v) IQ: support integer quantization. (vi) NN: support neural network quantization. (vii) STE: support Straight-Through Estimator, which permits gradients to propagate through during the backward pass that enables the continuation of the backpropagation algorithm.
Low-precision arithmetic has emerged as a pivotal technique for optimizing computational efficiency in scientific computing and machine learning applications, where reduced precision can significantly lower resource demands while maintaining acceptable accuracy (Higham and Pranesh, 2019). Several software libraries have been developed to simulate low-precision arithmetic, each with distinct capabilities tailored to specific use cases. In this section, we discuss the software packages listed in Table $2$The table’s design follows (Fasi and Mikaitis, 2023), highlighting their strengths and limitations in the context of low-precision arithmetic simulation.
GNU MPFR (Fousse et al., 2007) is a C library designed for multiple-precision floating-point computations with guaranteed correct rounding. It excels in scenarios requiring arbitrary precision for both the significand ($p$) and exponent range ($e$), making it a preferred choice for applications demanding high numerical accuracy, such as symbolic computation and numerical analysis. It supports rounding modes including round-to-nearest with ties-to-even (RNE), round-toward-zero (RZ), and round up/down (RUD). However, its default lack of subnormal number support ($s$, denoted as O) requires explicit user configuration, which can complicate workflows. Additionally, GNU MPFR does not offer built-in floating-point formats and is not suited for neural network training, limiting its applicability in machine learning contexts.
SIPE (Lefèvre, 2013), another C-based library, focuses on very low-precision computations with correct rounding. It supports multiple storage formats, restricted significand precision ($p$), and a restricted exponent range ($e$), while fully supporting subnormal numbers ($s$). SIPE implements RNE and RZ rounding modes, ensuring numerical stability for low-precision simulations, such as those in embedded systems. Its primary limitation lies in its restricted rounding mode support and lack of built-in formats, which reduces its versatility. Moreover, SIPE is not designed for neural network training, making it more suitable for general numerical simulations rather than machine learning applications.
rpe (Dawson and Düben, 2017), implemented in Fortran, is tailored for emulating reduced floating-point precision in large-scale numerical simulations, such as climate modeling. It operates with restricted significand ($p$) and exponent range ($e$) limited to built-in types (B) and supports subnormal numbers only for built-in types (B). A key advantage is its native support for the fp16 format, aligning with low-precision hardware standards like IEEE 754 binary16. However, rpe’s exclusive support for RNE rounding limits its flexibility in scenarios requiring diverse rounding strategies, and it does not support neural network training, focusing solely on numerical simulations.
FloatX (Flegar et al., 2019) and FlexFloat (Tagliavini et al., 2020), both C++ libraries, provide frameworks for customized floating-point arithmetic in low-precision simulations. They support restricted significand ($p$) and exponent range ($e$), fully support subnormal numbers ($s$), and use fp32/fp64 storage formats for compatibility with standard representations. Their simplicity, with support limited to RNE rounding, makes them accessible for educational purposes and prototyping. However, this restricted rounding support hampers their adaptability, and neither library includes built-in formats or supports neural network training, confining their use to general-purpose low-precision arithmetic experimentation. FloatX and FlexFloat is constrained by adhering strictly to the ”natural” C++ rules, such as the ”round-to-nearest, ties-to-even” rounding mode and standard datatype casting conventions. While these rules ensure consistency with native floating-point behavior, they may limit users who wish to explore a broader range of rounding strategies across various numerical simulation domains. This restriction can be a drawback for researchers and practitioners seeking more flexibility in modeling custom rounding behaviors tailored to specific applications.
INTLAB (Rump, 1999), a MATLAB toolbox, leverages interval arithmetic to facilitate low-precision floating-point simulation. It uses fp64 storage, supports restricted significand ($p$) and exponent range ($e$), and fully supports subnormal numbers ($s$). INTLAB provides RNE, RZ, and RUD rounding modes, offering moderate flexibility for numerical computations in MATLAB environments, such as verified computing. Its lack of built-in formats and optimization for neural network training limits its scope, positioning it as a tool for reliable numerical analysis rather than machine learning.
chop (Higham and Pranesh, 2019), another MATLAB library, enables low-precision arithmetic simulation with fp32/fp64 storage, restricted significand ($p$), and exponent range ($e$), alongside flexible subnormal number support (F). It supports built-in fp16 and bf16 formats, aligning with modern hardware standards, and implements RNE, RZ, RUD, and stochastic rounding (SR). The inclusion of SR is particularly valuable for simulating quantization effects, which are critical in machine learning research. Despite this, chop does not directly support neural network training and inference, and it is limited to the MATLAB environment, restricting its application to numerical experimentation and analysis of quantization impacts in machine learning.
CPFloat (Fasi and Mikaitis, 2023), a C library, is optimized for efficient low-precision arithmetic simulation, supporting fp32/fp64 storage, restricted significand ($p$), and exponent range ($e$), with flexible subnormal number support (F). It includes built-in fp16, bf16, and tf32 formats, enhancing compatibility with hardware-accelerated systems such as GPUs. CPFloat’s comprehensive support for rounding modes—RNE, RNZ, RNA, RZ, RUD, RO, and SR—makes it highly versatile for numerical simulations requiring diverse rounding strategies. However, CPFloat is not designed for neural network training, focusing instead on general-purpose low-precision arithmetic, such as in algorithm development and hardware emulation.
QPyTorch (Zhang et al., 2019) is specifically designed for low-precision arithmetic for neural network training. QPyTorch, a PyTorch-integrated Python tool, uses fp32 storage with restricted $p$ and $e$, lacks $s$ support, and offers RNE, RNZ, and SR rounding modes, enabling efficient training (e.g., CIFAR10) via fused kernels. However, its limited rounding support (lacking RZ, RUD, RO) restricts advanced quantization studies and numerical simulations.
All aforementioned libraries, except QPyTorch, lack Straight-Through Estimator (STE) support and are thus restricted to post-training quantization (PTQ), rendering them ineffective for quantization-aware training (QAT). Our Pychop library supports fp32/FP64 storage and built-in fp16 and bf16 formats, constrained precision ($p$) and exponent ($e$), and flexible significand ($s$), with rounding modes (RNE, RZ, RUD, RO, SR). It automatically detects tensor gradient information to enable STE, enhancing quantization research flexibility. Pychop integrates seamlessly with NumPy, JAX, and PyTorch—outperforming QPyTorch’s PyTorch-only scope—and offers efficient rounding. Its versatility and multi-framework compatibility make Pychop a superior tool for quantized neural network training across diverse deep learning workflows and scientific computations.
3. The Pychop library
Real numbers in digital systems must be approximated using discrete representations due to finite storage constraints. Two fundamental approaches dominate this domain: floating-point representation, which excels in representing a wide range of values with variable precision, and fixed-point representation, which prioritizes simplicity and efficiency in constrained environments. This section provides rigorous definitions for both methods, with a particular focus on the IEEE 754 standard for floating-point representation, and compares their theoretical and practical implications in computational tasks. Our mixed-precision emulation software is implemented as a modular Python code within NumPy, PyTorch, and JAX, comprising three primary components: the Chop/Chopf/Chopi class for core quantization logic and various components in the layers class for neural network integration, supporting various roundings (see Table 3 for details) and subnormal numbers. Here, we detail its design and implementation, emphasizing its main principles and user-friendly features.
Table 3. Rounding Modes for the rmode in low precision emulation.
rmode Rounding Mode Description
1 Round to nearest, ties to even Rounds to the nearest representable value; in cases of equidistance, selects the value with an even least significant digit (IEEE 754 standard).
2 Round toward $\infty$ (round up) Rounds to the smallest representable value greater than or equal to the input, directing towards positive infinity.
3 Round toward $\infty$- ∞ (round down) Rounds to the largest representable value less than or equal to the input, directing towards negative infinity.
4 Round toward zero Discards the fractional component, yielding the integer closest to zero without exceeding the input’s magnitude.
5 Stochastic (proportional) Employs probabilistic rounding where the probability of rounding up is proportional to the fractional component.
6 Stochastic (uniform) Applies probabilistic rounding with an equal probability (0.5) of rounding up or down, independent of the fractional value.
7 Round to nearest, ties to zero Rounds to the nearest representable value; in cases of equidistance, selects the value closer to zero.
8 Round to nearest, ties away Rounds to the nearest representable value; in cases of equidistance, selects the value farther from zero.
9 Round toward odd Rounds to the nearest representable odd value; in cases of equidistance, selects the odd value in the direction of the original number.
3.1. Floating-point arithmetic
Floating-point representation approximates real numbers using a binary format analogous to scientific notation. Pychop emulates floating-point arithmetic by decomposing a tensor into sign, exponent, and significand components, following IEEE 754 conventions; The IEEE 754 standard, established in 1985 and revised in 2008 and 2019, provides a widely adopted framework for floating-point arithmetic, ensuring consistency across hardware and software implementations.
A floating-point number $x$ in the IEEE 754 standard is defined based on its encoding as a tuple of three components: a sign bit, an exponent, and a mantissa (or significand). Mathematically, the value $x$ is expressed as
$x=(-1)^{s}\cdot M\cdot 2^{E},$
where $s\in\{0,1\}$ is the sign bit ($s=0$ for positive, $s=1$ for negative), $M$ is the mantissa, a binary fraction interpreted based on the exponent field, and $E$ is the exponent—an integer derived from the stored exponent field with a bias adjustment..
The IEEE 754 standard defines several formats, with the most common being single precision (32 bits) and double precision (64 bits). For a format with a $k$-bit exponent and a $p-1$-bit mantissa fraction, the bit layout is
$[s\,|\,e\,|\,m],$
where $s$ is the 1-bit sign, $e$ is the $k$-bit exponent field, and $m$ is the $p-1$-bit fractional part of the mantissa—the total precision $p$ includes an implicit leading bit for normalized numbers.
The interpretation of $M$ and $E$ depends on the value of the exponent field $e$:
  • Normalized Numbers ($21\leq e\leq 2^{k}-$ 2): The mantissa is $M=1+\sum_{i=1}^{p-1}m_{i}\cdot 2^{-i}$, where $m_{i}$ are the bits of the fractional field, and the exponent is $E=e-\text{bias}$, with $\text{bias}=2^{k-1}-1$. Thus
    $x=(-1)^{s}\cdot\left(1+\sum_{i=1}^{p-1}m_{i}\cdot 2^{-i}\right)\cdot 2^{e-\text{bias}}.$
  • Denormalized Numbers ($e=0$): The mantissa is $M=0+\sum_{i=1}^{p-1}m_{i}\cdot 2^{-i}$ (no implicit leading 1), and the exponent is $E=1-\text{bias}$. Thus
    $x=(-1)^{s}\cdot\left(\sum_{i=1}^{p-1}m_{i}\cdot 2^{-i}\right)\cdot 2^{1-\text{bias}}.$
    Denormalized numbers allow representation of values closer to zero, mitigating the abrupt underflow of normalized numbers.
  • Special Values:
    • If $e=2^{k}-1$ and $m=0$, then $x=(-1)^{s}\cdot\infty$ (infinity).
    • If $e=2^{k}-1$ and $m\neq 0$, then $x$ represents a Not-a-Number (NaN), used to indicate invalid operations.
    • If $e=0$ and $m=0$, then $x=(-1)^{s}\cdot 0$ (signed zero).
3.2. Implementation details
Floating-point arithmetic often produces results that cannot be represented exactly within the finite precision $p$. The IEEE 754 standard defines several rounding modes to map an exact real number $r\in\mathbb{R}$ to a representable floating-point number $x\in\mathbb{F}$. Let $\lfloor r\rfloor_{\mathbb{F}}$ and $\lceil r\rceil_{\mathbb{F}}$ denote the closest representable numbers in $\mathbb{F}$ such that $\lfloor r\rfloor_{\mathbb{F}}\leq r\leq\lceil r\rceil_{\mathbb{F}}$. The midpoint between two consecutive representable numbers is $m=\frac{\lfloor r\rfloor_{\mathbb{F}}+\lceil r\rceil_{\mathbb{F}}}{2}$.
To round to low precision binary floating point numbers, Pychop provides two modules, namely Chop and LightChop. Chop features a greater set of functionalities and follows the design of MATLAB chop; LightChop is deigned to be a lightweight “Chop” with essential features and enables efficient vectorization. Different backends of LightChop offer different features. The Torch and JAX backend offers GPU deployment, while the NumPy backend leverages Dask (Rocklin, 2015), a parallel computing library, to process a large NumPy array by chunking the array with user-defined chunk size, where the chunk size dictates how the array is split into smaller, manageable chunks for parallel or out-of-core computation, enabling efficient, scalable processing through lazy evaluation and a task graph that executes only when triggered, balancing memory use and parallelism based on the chosen chunk size.
The calling of Pychop is straightforward. In the following, we will illustrate how to use Chop and LightChop to emulate low precision arithmetic. The prototype of Chop and LightChop are separately described as below:
  • 1Chop(prec:=str=’h’, subnormal:=bool=True, rmode:=int=1, flip:=bool=False, explim:=int=1, p: =float=0.5, randfunc=None, customs:Customs=None, random_state:=int=0, verbose:=int=0)
    This interface is designed to support detailed control over precision, range, and rounding behavior in floating-point operations, allowing users to specify the target arithmetic precision (prec, defaulting to ’h’; if customs is fed, this parameter will be omitted), whether subnormal numbers are supported (subnormal, defaulting to True), and the rounding mode (rmode, with options like “nearest” or stochastic methods, defaulting to 1). Additional features include an option to randomly flip bits in the significand for error simulation (if flip, which defaults to False, is set to True, then each element of the rounded result has, with probability p (default 0.5), a randomly chosen bit in its significand flipped), to control exponent limits (explim, defaulting to True), and a custom random function for stochastic rounding (randfunc, defaulting to None). Users can also define custom precision parameters via a dataclass Customs(emax, t, exp_bits, sig_bits) where emax refers to the maximum value of the exponent, t refers to the significand bits which includes the hidden bit, and exp_bits and sig_bits refer to the exponent bit and significand bit which excludes the hidden bit, respectively. random_state is used to set a random seed for reproducibility, and toggle verbosity for unit-roundoff output (verbose, defaulting to 0).
  • 1LightChop(exp_bits:=int, sig_bits:=int, chunk_size: =int=1000, rmode:=int=1, subnormal:=bool=True, chunk_size: =int=1000, random_state:=int=42)
    This interface facilitates precise control over the precision range and rounding behavior of floating-point operations. It enables users to specify the bitwidth for the exponent (exp_bits) and significand (sig_bits) of floating-point numbers. A rounding mode parameter (rmode), defaulting to 1, is included to govern rounding behavior. Subnormal numbers are managed via the subnormal parameter, which defaults to True, i.e., all subnormal numbers are inherently supported. The chunk_size parameter defines the number of elements within each smaller sub-array (or chunk) into which a large array is segmented for parallel processing. Smaller chunk sizes enhance parallelism but incur greater overhead, whereas larger chunks minimize overhead at the expense of increased memory requirements. In essence, chunk_size serves as the fundamental unit of work managed by Dask, striking a balance between computational efficiency and memory limitations. This parameter is exclusively applicable when using the NumPy backend. Additionally, a random seed (random_state), defaulting to 0, can be configured to ensure reproducibility in stochastic rounding scenarios.
Generally, LightChop is faster than Chop, which will be verified in Section 4, but Chop includes more floating point emulation features. We demonstrate their basic usage below:
1from Pychop import Chop
2import numpy as np
3
4X = np.random.randn(5000, 5000)
5ch = Chop(’h’, rmode=1) # Standard IEEE 754 half precision
6# other parameters are left as default.
7Xq = ch(X) # Rounding values
Use built-in precision
1import numpy as np
2from Pychop import Chop, LightChop
3from Pychop import Customs
4
5X = np.random.randn(5000, 5000)
6ch = Chop(customs=Customs(exp_bits=5, sig_bits=10), rmode=1) # half precision (5 exponent bits, 10+(1) significand bits, (1) is implicit bits)
7Xq = ch(X)
8
9ch = LightChop(exp_bits=5, sig_bits=10, rmode=1)
10Xq = ch(X)
Customized precision
3.3. Fixed-point arithmetic
Fixed-point is associated with numbers of a fixed number of bits, allocating a predetermined number of bits to the integer part and the fractional part, unlike floating-point representations that use an exponent and significand. The shift to fixed-point arithmetic is driven by several key advantages. First, fixed-point compute units are typically faster and significantly more efficient in terms of hardware resources and power consumption compared to floating-point units. Their smaller logic footprint allows for a higher density of compute units within a given area and power budget. Second, low-precision data representation reduces the memory footprint, enabling larger models to fit within available memory while also lowering bandwidth requirements. Fixed-point operations align well with digital signal processors (DSPs) and field-programmable gate arrays (FPGAs) because many lack dedicated floating-point units or optimize for fixed-point arithmetic. Collectively, these benefits enhance data-level parallelism, leading to substantial performance gains (Gupta et al., 2015).
Fixed-point representation provides a simpler alternative to floating-point by fixing the position of the binary point within a binary number. This conversion from floating point numbers to fixed-point numbers employs a fixed scaling factor (implicit or explicit), enabling fractional values to be represented as integers scaled by a constant, such as $2^{-f}$, where $f$ is the number of fractional bits.
A fixed-point number $x$ is defined as an integer $I$ scaled by a fixed factor $2^{-f}$, where $f$ is the number of fractional bits:
$x=I\cdot 2^{-f}$
where $I$ is an integer, which may be signed (typically in two’s complement) or unsigned, $f$ is the fixed number of fractional bits, and $n$ is the total number of bits is, with $n-f$ bits allocated to the integer part and $f$ bits to the fractional part.
The binary representation of $I$ can be written as:
$I=b_{n-1}b_{n-2}\ldots b_{f}\cdot b_{f-1}\ldots b_{0}$
with the binary point implicitly placed between bits $b_{f}$ and $b_{f-1}$. The numerical value is
$x=\left(\sum_{i=0}^{n-1}b_{i}\cdot 2^{i}\right)\cdot 2^{-f}.$
For a signed representation using two’s complement, the most significant bit $b_{n-1}$ is the sign bit, and the value of $I$ is
$I=-b_{n-1}\cdot 2^{n-1}+\sum_{i=0}^{n-2}b_{i}\cdot 2^{i}.$
Thus, the fixed-point value $x$ becomes
$x=\left(-b_{n-1}\cdot 2^{n-1}+\sum_{i=0}^{n-2}b_{i}\cdot 2^{i}\right)\cdot 2^{-f}.$
For an unsigned representation, the sign bit is absent, and $I=\sum_{i=0}^{n-1}b_{i}\cdot 2^{i}$, so
$x=\left(\sum_{i=0}^{n-1}b_{i}\cdot 2^{i}\right)\cdot 2^{-f}.$
3.3.1. Process of fixed-point quantization
The process of fixed-point quantization in neural networks involves several steps. A fixed-point number is typically denoted as $Qm.f$, where $m$ represents the number of integer bits and $f$ the number of fractional bits. For instance, a $Q8.8$ format uses 8 bits for the integer part and 8 bits for the fractional part, stored as a 16-bit integer. To convert a floating-point value $x$ to a fixed-point value $q$, the value is scaled and rounded according to
$q=\text{round}(x\cdot 2^{f}).$
The prototype of Pychop for fixed-point quantization is as below:
1Chopf(ibits:=int=4, fbits:=int=4, rmode:=int=1)
where the ibits refers to the bitwidth of integer part, fbits refers to the bitwidth of fractional part, and rmode indicates the rounding mode used in (3.3.1); the supported rounding modes can be found in Table 3.
The usage is given by the following example.
1from Pychop import Chopf
2import numpy as np
3
4X = np.random.randn(5000, 5000)
5ch = Chopf(ibits=4, fbits=4, rmode=1)
6ch(X)
Fix-point representation
3.3.2. Relation to and Deep Learning Deployment
The adoption of fixed-point quantization in deep learning offers several benefits and significant contributions that enhance its utility across various applications, and plays a pivotal role in neural networks by optimizing both inference and training phases. By leveraging fixed-point arithmetic, which uses fewer bits—such as 16-bit $Q8.8$ representations compared to 32-bit floats—this technique substantially reduces the memory footprint and accelerates multiply-accumulate (MAC) operations, which are fundamental to DNN computation. Its compatibility with hardware is a key advantage, as many edge devices like microcontrollers natively support fixed-point operations, eliminating the need for floating-point emulation and thereby boosting performance. With careful selection of the integer ($m$) and fractional ($f$) bit allocations, fixed-point quantization maintains accuracy close to that of floating-point models, a capability validated by research (Lin et al., 2016b). Ultimately, this technique has enabled the deployment of sophisticated DNNs on resource-limited platforms, significantly broadening the practical impact of AI in fields like mobile computing, real-time image processing, and autonomous navigation.
In the context of neural networks and deep learning, fixed-point quantization is applied to weights, activations, and sometimes gradients to optimize computation and storage, making it a cornerstone for efficient deployment and training. For inference, it replaces costly floating-point operations with fixed-point arithmetic, which is natively supported by many embedded systems. In training, quantization-aware training ensures the model adapts to reduced precision, minimizing accuracy loss. Fully fixed-point training, though less common, has been studied for end-to-end optimization on fixed-point hardware. This compatibility with hardware acceleration—particularly DSPs and FPGAs, which often lack native floating-point units or optimize for fixed-point operations—bridges the gap between complex DNNs (e.g., convolutional neural networks or Transformers) and practical, low-power deployment. Fixed-point quantization thus reduces memory and computational costs while enabling practical deployment of DNNs on edge devices, as explored in foundational works like (Lin et al., 2016a).
3.4. Integer arithmetic
Integer quantization is a fundamental technique in digital systems to approximate real numbers using a finite set of integers, enabling efficient storage and computation in applications such as machine learning, digital signal processing, and embedded systems. Two primary approaches to quantization exist: symmetric quantization, which balances positive and negative ranges around zero, and asymmetric quantization, which allows unequal ranges for positive and negative values. In principle, integer quantization maps a real number $r\in\mathbb{R}$ to a discrete integer value $x\in\mathbb{Z}$ through scaling and rounding. The process can be adapted for either symmetric or asymmetric quantization depending on the range and scaling strategy. Consider a real number $r$ within a specified range $[r_{\text{min}},r_{\text{max}}]$. The goal is to represent $r$ using $n$-bit integers, defining a discrete set of quantization levels. The general quantization process involves scaling, rounding, and clamping, with variations depending on whether symmetric or asymmetric quantization is used.
In the following, we briefly explain the symmetric quantization and asymmetric quantization.
  • Symmetric Quantization: Symmetric quantization balances the quantization range around zero, ensuring that the positive and negative ranges are equal in magnitude. This is particularly useful in applications where data distributions (e.g., neural network weights) are centered around zero.
    For an $n$-bit signed integer in two’s complement, the representable range is $[-2^{n-1},2^{n-1}-1]$. The quantization range is defined symmetrically as $[-\omega,\omega]$, where $\omega$ is the maximum absolute value to be represented. The scaling factor $\Delta$ is:
    $\Delta=\frac{\omega}{2^{n-1}-1}.$
    The scaled value $s$ is computed as:
    $s=\frac{r}{\Delta}.$
    The integer $x$ is obtained by rounding:
    $\overline{x}=\mathrm{R}(s),\quad\text{ where }\mathrm{R}(\cdot)\text{ denotes general rounding operator.}$.
    Particularly, if $\mathrm{R}$ is round to nearest, ties to even, i.e., $\mathrm{R}\equiv\mathrm{RNE}$, then $\mathrm{R}(s)=\lfloor s+0.5\rfloor$.
    Sometimes, a clamping is applied when needed, i.e.,
    $x=\mathrm{max}(-2^{n-1},\mathrm{min}(\overline{x},2^{n-1}-1)).$
    To map $x$ back to $\hat{r}$, the dequantization is applied:
    $\hat{r}=x\cdot\Delta.$
    Here, $r_{\text{min}}=-\omega$ and $r_{\text{max}}=\omega$, ensuring symmetry around zero. When $r=0$, the quantized value is $x=0$, preserving symmetry without an offset.
  • Asymmetric Quantization: Asymmetric quantization allows unequal ranges for positive and negative values, typically by defining a range $[r_{\text{min}},r_{\text{max}}]$ where $r_{\text{min}}\neq-r_{\text{max}}$. This is useful when the data distribution is skewed (e.g., all positive values in rectified activations like ReLU).
    For an $n$-bit signed integer, the range is still $[-2^{n-1},2^{n-1}-1]$, but the quantization maps $[r_{\text{min}},r_{\text{max}}]$ to this range. The scaling factor $\Delta$ is
    $\Delta=\frac{r_{\text{max}}-r_{\text{min}}}{2^{n}-1}.$
    The scaled value $s$ is
    $s=\frac{r-r_{\text{min}}}{\Delta}.$
    Similarly, the integer $x$ is obtained by rounding and clamping:
    $\overline{x}$ $=\mathrm{R}(s),\quad\text{ where }\mathrm{R}(\cdot)\text{ denotesgeneral rounding operator,}$,
    $x$ $=\mathrm{max}(-2^{n-1},\mathrm{min}(\overline{x},2^{n-1}-1)).$
    Dequantization maps $x$ back to $\hat{r}$:
    $\hat{r}=r_{\text{min}}+x\cdot\Delta.$
    In asymmetric quantization, the zero point (where $r=0$) maps to an integer $z$ in the quantized domain:
    $z=\mathrm{R}\left(\frac{0-r_{\text{min}}}{\Delta}\right).$
    This introduces an offset, which may require additional computation during arithmetic operations.
Integer quantization also includes uniform quantization and non-uniform quantization. Uniform quantization refers to dividing an integer range into equally-sized segments, while non-uniform quantization refers to using segments of varying sizes, often adjusted based on the integer values’ distribution or importance. Compared to non-uniform quantization, uniform quantization is computationally efficient, hardware-friendly, and easier to implement. Most modern processors, including CPUs, GPUs, and specialized accelerators like TPUs and NPUs, are optimized for integer arithmetic with uniform quantization, enabling fast matrix multiplications and reduced memory overhead. While non-uniform quantization can provide better precision for highly skewed data distributions, it often requires more complex lookup tables or clustering methods, which increase computational cost and slow down inference. As a result, uniform quantization remains the standard choice for deep learning acceleration in both training and inference. Therefore, Pychop focuses on uniform quantization.
In the following, we demonstrate the usage of Pychop for integer quantization:
  • 1class Chopi(bits=8, symmetric=False, per_channel=False, axis=0)
    The Chopi framework offers tailored integer quantization, allowing users to specify the bitwidth length (bits) and choose between symmetric or asymmetric quantization (symmetric). Additionally, two channel-specific parameters enable further customization: per_channel determines whether quantization is applied on a per-channel basis, while axis specifies the dimension along which channel-wise quantization occurs when is set to True. Below is a simple demonstration:
    1from Pychop import Chopi
    2
    3ch = Pychop.Chopi(bits=8, symmetric=False, per_channel=False, axis=0)
    4X_q = ch.quantize(X_np) # to integers
    5X_inv = ch.dequantize(X_q) # back to floating point numbers
3.5. Common mathematical functions support and array manipulation routines
We simulate common mathematical functions and operations (such as built-in functions in NumPy, PyTorch, or JAX) in low-precision arithmetic by first rounding the input to low precision, performing operations in the working precision, and then rounding the result back to low precision. This approach contrasts with CPFloat, which applies mathematical operations in working precision to inputs in working precision before rounding the final result to low precision.
Table 4. Functions Support and Array Manipulation Routines (Part 1)
33footnotemark: 3
Function Description
Trigonometric Functions
sin Computes sine. Input in radians; output in $[-1,1]$.
cos Computes cosine. Input in radians; output in $[-1,1]$.
tan Computes tangent. Input in radians; discontinuities at $\pi/2+k\pi$.
arcsin Computes arcsin. Input in $[-1,1]$; output in $[-\pi/2,\pi/2]$ radians.
arccos Computes arccos. Input in $[-1,1]$; output in $[0,\pi]$ radians.
arctan Computes arctan. Output in $[-\pi/2,\pi/2]$ radians.
Hyperbolic Functions
sinh Computes hyperbolic sine. Output unrestricted.
cosh Computes hyperbolic cosine. Output non-negative.
tanh Computes hyperbolic tangent. Output in $(-1,1)$.
arcsinh Computes inverse hyperbolic sine. Output in real numbers.
arccosh Computes inverse hyperbolic cosine. Input $\geq 1$; output in $[0,\infty)$.
arctanh Computes inverse hyperbolic tangent. Input in $(-1,1)$; output real.
Exponential and Logarithmic Functions
exp Computes $e^{x}$. Input unrestricted; output positive.
expm1 Computes $e^{x}-1$. Enhanced precision for small $x$.
log Computes natural logarithm (base $e$). Input positive; output unrestricted.
log10 Computes base-10 logarithm. Input positive; output unrestricted.
log2 Computes base-2 logarithm. Input positive; output unrestricted.
log1p Computes $\log(1+x)$. Input $>-1$; enhanced precision for small $x$.
Power and Root Functions
sqrt Computes square root. Input non-negative; output non-negative.
cbrt Computes cube root. Input unrestricted; output sign matches input.
Aggregation and Linear Algebra Functions
sum Computes sum of array elements along axis.
prod Computes product of array elements along axis.
mean Computes mean of array elements along axis.
std Computes standard deviation of array elements along axis.
var Computes variance of array elements along axis.
dot Computes dot product of two arrays.
matmul Computes matrix multiplication of two arrays.
Special Functions
erf Computes error function. Output in $(-1,1)$.
erfc Computes complementary error function ($1-\text{erf}$).
gamma Computes gamma function. Input unrestricted.
Other Mathematical Functions
fabs Computes floating-point absolute value. Output non-negative.
logaddexp Computes logarithm of sum of exponentials.
cumsum Computes cumulative sum along axis.
cumprod Computes cumulative product along axis.
degrees Converts radians to degrees.
radians Converts degrees to radians.
Table 5. Functions Support and Array Manipulation Routines (Part 2)
 [^4] All functions are computed with chopping to enforce low-precision format, where applicable.
Function Description
Rounding and Clipping Functions
floor Computes floor. Rounds down to nearest integer.
ceil Computes ceiling. Rounds up to nearest integer.
round Rounds to specified decimals.
sign Computes sign. Returns $-1$, 00, or $1$.
clip Clips values to range $[a_{\text{min}},a_{\text{max}}]$.
Miscellaneous Functions
abs Computes absolute value. Output non-negative.
reciprocal Computes $1/x$. Input must not be zero.
square Computes square of input. Output non-negative.
Additional Mathematical Functions
frexp Decomposes into significand and exponent. Chopping on significand.
hypot Computes $\sqrt{x^{2}+y^{2}}$. Inputs are real numbers.
diff Computes difference between consecutive array elements.
power Computes element-wise $x^{y}$.
modf Decomposes into fractional and integral parts. Chopping on fractional part.
ldexp Multiplies by $2$ to exponent power.
angle Computes phase angle of complex number. Output in radians.
real Extracts real part of complex number.
imag Extracts imaginary part of complex number.
conj Computes complex conjugate.
maximum Computes element-wise maximum of two inputs.
minimum Computes element-wise minimum of two inputs.
Binary Arithmetic Functions
multiply Computes element-wise product.
mod Computes element-wise modulo. Divisor must not be zero.
divide Computes element-wise division. Divisor must not be zero.
add Computes element-wise sum.
subtract Computes element-wise difference.
floor_divide Computes element-wise floor division. Divisor must not be zero.
bitwise_and Computes bitwise AND of integer inputs.
bitwise_or Computes bitwise OR of integer inputs.
bitwise_xor Computes bitwise XOR of integer inputs.
The usage of common functions is illustrate as an example below:
1ch = Chop(’q43’, rmode=1)
2ch.sin(X)
3.6. Seamless PyTorch / JAX Integration
Pychop currently supports NumPy’s array, PyTorch’s tensor, and JAX’s array as input for their respective computing routines. Utilizing PyTorch, e.g., enables faster vectorized computation, making it efficient for large-scale datasets.
In the example below, we generate three distinct inputs: a NumPy array, a PyTorch tensor, and a JAX array. By specifying the appropriate backend, we can configure Pychop to use 5 exponent bits and 10 significand bits with round to nearest mode to process these inputs accordingly. For instance, selecting the “torch” backend allows Pychop to handle X_th, the PyTorch tensor. Additionally, GPU deployment can be enabled by transferring the PyTorch tensor or JAX array to a GPU device, such as with X_th.to(’cuda’).
1import torch, pychop, jax
2from pychop import LightChop
3
4X_np = np.random.randn(5000, 5000) # Numpy array
5X_th = torch.Tensor(X_np) # torch array
6X_jx = jax.numpy.asarray(X_np)
7
8pychop.backend(’numpy’) # Use NumPy backend. NumPy is the default option.
9ch = LightChop(exp_bits=5, sig_bits=10, rmode=1)
10emulated= ch(X_np)
11
12pychop.backend(’torch’) # Use PyTorch backend.
13ch = LightChop(exp_bits=5, sig_bits=10, rmode=1)
14emulated= ch(X_th)
15
16pychop.backend(’jax’) # Use JAX backend.
17ch = LightChop(exp_bits=5, sig_bits=10, rmode=1)
18emulated= ch(X_jx)
3.7. Neural network quantization settings
Mixed-precision Deep Neural Networks provide the energy efficiency and throughput essential for hardware deployment, particularly in resource-limited settings, often without compromising accuracy. However, identifying the optimal per-layer bit precision remains challenging due to the vast search space introduced by the diverse range of models, datasets, and quantization techniques (see (Rakka et al., 2024) and references therein). Neural network training and inference are inherently resilient to errors, a characteristic that distinguishes them from traditional workloads that demand precise computations and high dynamic range number representations. It is well understood that, given the presence of statistical approximation and estimation errors, high-precision computations in learning tasks are often unnecessary (Bottou and Bousquet, 2007). Furthermore, introducing noise during training has been shown to improve neural network performance (Bishop, 1995; Audhkhasi et al., 2013; Kosko et al., 2020).
Pychop is well-suited for post-quantization and quantization-aware training for neural network deployment, including quantization-aware training (QAT) and post-training quantization (PTQ). Its design prioritizes simplicity and flexibility, making it an ideal tool for experimenting with and fine-tuning quantization strategies. In the following, we provide a concise illustration of how Pychop can be effectively utilized in quantization applications for neural networks, demonstrating its ease of use and integration into existing workflows. This process is adapted as follows:
  • Training: During quantization-aware training (QAT), the network simulates fixed-point arithmetic by quantizing weights and activations in the forward pass. Gradients may remain in higher precision.
  • Inference: Weights and activations are quantized to required format for efficient computation.
3.7.1. Principle and basic usage
Table 6. Common implemented quantized Layers (part) and their original PyTorch names
Quantized Layer Name Original PyTorch Name
QuantizedLinear / FPQuantizedLinear / IntQuantizedLinear nn.Linear
QuantizedConv1d / FPQuantizedConv1d / IntQuantizedConv1d nn.Conv1d
QuantizedConv2d / FPQuantizedConv2d / IntQuantizedConv2d nn.Conv2d
QuantizedConv3d / FPQuantizedConv3d / IntQuantizedConv3d nn.Conv3d
QuantizedRNN / FPQuantizedRNN / IntQuantizedRNN nn.RNN
QuantizedLSTM / FPQuantizedLSTM / IntQuantizedLSTM nn.LSTM
QuantizedMaxPool1d / FPQuantizedMaxPool1d / IntQuantizedMaxPool1d nn.MaxPool1d
QuantizedMaxPool2d / FPQuantizedMaxPool2d / IntQuantizedMaxPool2d nn.MaxPool2d
QuantizedMaxPool3d / FPQuantizedMaxPool3d / IntQuantizedMaxPool3d nn.MaxPool3d
QuantizedAvgPool / FPQuantizedAvgPool / IntQuantizedAvgPool nn.AvgPool
QuantizedAttention / FPQuantizedAttention / IntQuantizedAttention nn.Attention
QuantizedBatchNorm1d / FPQuantizedBatchNorm1d / IntQuantizedBatchNorm1d nn.BatchNorm1d
QuantizedBatchNorm2d / FPQuantizedBatchNorm2d / IntQuantizedBatchNorm2d nn.BatchNorm2d
QuantizedBatchNorm3d / FPQuantizedBatchNorm3d / IntQuantizedBatchNorm3d nn.BatchNorm3d
Table 7. Quantized Optimizers (part) and Their Original PyTorch Names
Common quantized optimizer name Original PyTorch name
QuantizedSGD / FPQuantizedSGD / IntQuantizedSGD torch.optim.SGD
QuantizedAdam / FPQuantizedAdam / IntQuantizedAdam torch.optim.Adam
QuantizedRMSProp / FPQuantizedRMSProp / IntQuantizedRMSProp torch.optim.RMSprop
QuantizedAdagrad / FPQuantizedAdagrad / IntQuantizedAdagrad torch.optim.Adagrad
QuantizedAdadelta / FPQuantizedAdadelta / IntQuantizedAdadelta torch.optim.Adadelta
QuantizedAdamW / FPQuantizedAdamW / IntQuantizedAdamW torch.optim.AdamW
Pychop simulates multiple-precision neural network training by introducing floating-point / fixed-point / integer quantization into the training process while still performing the underlying computations in full precision (e.g., fp32/FP64) using PyTorch’s native capabilities. The pre-built quantized layer and optimizers class extend the multiple precision emulation of torch.nn.Module and algorithms in torch.optim, applying the simulator to various layer and arithmetic operations. The part of the implemented quantized layers and optimization algorithms are listed in Table 6 and Table 7. All layers and optimizers follow a modular design for easy extension, with the same parameter settings of the original PyTorch modules with additional parameters to define rounding modes and quantization settings such as bitwidth for exponent and significand (exp_bits and sig_bits), and preserve original tensor shapes and PyTorch compatibility. As for optimizers, the quantization will be applied to gradients, momenta, and other state variables used by the optimizers. The design of this functionality facilitates the study of quantization effects in neural network performance, the simulation of low-precision hardware, and the evaluation of numerical stability in deep learning. We briefly summarize these functions as follows:
  • Implementation: For layers, Pychop quantizes weights, input, and bias before operations, then uses standard PyTorch matrix multiplication and addition with working precision (either fp32 or fp64, which depends on user settings). The gradient flow through the quantized operations is maintained in working precision.
  • Parameters: Pychop allows the quantization of weights and biases during initialization or forward pass, and quantizes inputs, performs matrix multiplication, and adds quantized bias, all in the specified format.
  • Flexibility: Pychop allows the quantization of different parts of the training process independently, such as weights, activations, gradients, momentum, and gradient accumulators. It also provides the pre-built layers and optimizers for training. It supports customizable low-precision formats, including floating-point (with configurable bitwidth for exponent and significand parts), fixed-point (with configurable bitwidth for integer and fraction parts), and integers arithmetic (with configurable bitwidth for integer part).
  • Extensibility: Template design supports adaptation to convolutional or recurrent layers. It also provides built-in quantized layers, for example, QuantizedLinear, QuantizedRNN, QuantizedLSTM, QuantizedGRU, which corresponds to the low precision emulation of nn.Linear, nn.RNN, nn.LSTM, and nn.GRU.
In the following, we demonstrate how to use the derived layer modules in Pychop.layers (following Table 6) to build quantization-aware training for convolutional neural networks.
1class CNN(nn.Module):
2 def =__init__(self):
3 =super().=__init__()
4 self.conv1 = nn.Conv2d(1, 16, 3, 1, 1)
5 self.pool = nn.MaxPool2d(2, 2)
6 self.conv2 = nn.Conv2d(16, 32, 3, 1, 1)
7 self.fc1 = nn.Linear(32 * 7 * 7, 128)
8 self.fc2 = nn.Linear(128, 10)
9
10 def =forward(self, x):
11 x = F.relu(self.conv1(x))
12 x = self.pool(x)
13 x = F.relu(self.conv2(x))
14 x = self.pool(x)
15 x = x.view(-1, 32 * 7 * 7)
16 x = F.relu(self.fc1(x))
17 x = self.fc2(x)
18 return x
Use built-in precision
1class QuantizedCNN(nn.Module):
2 def =__init__(self, exp_bits=5, sig_bits=10, rmode=1):
3 =super().=__init__()
4 self.conv1 = QuantizedConv2d(1, 16, 3, exp_bits, sig_bits, rmode=rmode)
5 self.pool = QuantizedMaxPool2d(2, exp_bits, sig_bits, rmode=rmode)
6 self.conv2 = QuantizedConv2d(16, 32, 3, exp_bits, sig_bits, rmode=rmode)
7 self.fc1 = QuantizedLinear(32 * 5 * 5, 128, exp_bits, sig_bits, rmode=rmode)
8 self.fc2 = QuantizedLinear(128, 10, exp_bits, sig_bits, rmode=rmode)
9
10 def =forward(self, x):
11 x = F.relu(self.conv1(x))
12 x = self.pool(x)
13 x = F.relu(self.conv2(x))
14 x = self.pool(x)
15 x = x.view(x.size(0), -1)
16 x = F.relu(self.fc1(x))
17 x = self.fc2(x)
18 return x
Specify floating point precision.
1class QuantizedCNN(nn.Module):
2 def =__init__(self):
3 =super().=__init__()
4 self.conv1 = FPQuantizedConv2d(1, 16, 3, 1, 1, ibits=8, fbits=8)
5 self.pool = FPQuantizedMaxPool2d(2, 2, ibits=8, fbits=8)
6 self.conv2 = FPQuantizedConv2d(16, 32, 3, 1, 1, ibits=8, fbits=8)
7 self.fc1 = FPQuantizedLinear(32 * 7 * 7, 128, ibits=8, fbits=8)
8 self.fc2 = FPQuantizedLinear(128, 10, ibits=8, fbits=8)
9
10 def =forward(self, x):
11 x = F.relu(self.conv1(x))
12 x = self.pool(x)
13 x = F.relu(self.conv2(x))
14 x = self.pool(x)
15 x = x.view(-1, 32 * 7 * 7)
16 x = F.relu(self.fc1(x))
17 x = self.fc2(x)
18 return x
Specify fixed-point precision.
1class QuantizedCNN(nn.Module):
2 def =__init__(self, bits=8):
3 =super(QuantizedCNN, self).=__init__()
4 self.conv1 = IntQuantizedConv2d(1, 16, 3, padding=1, bits=bits)
5 self.pool = nn.MaxPool2d(2, 2)
6 self.conv2 = IntQuantizedConv2d(16, 32, 3, padding=1, bits=bits)
7 self.fc1 = IntQuantizedLinear(32 * 7 * 7, 128, bits=bits)
8 self.fc2 = IntQuantizedLinear(128, 10, bits=bits)
9
10 def =forward(self, x):
11 x = F.relu(self.conv1(x))
12 x = self.pool(x)
13 x = F.relu(self.conv2(x))
14 x = self.pool(x)
15 x = x.view(-1, 32 * 7 * 7)
16 x = F.relu(self.fc1(x))
17 x = self.fc2(x)
18 return x
Specify bitwidth for integer.
For low precision optimizers, similarly, one can define the derived class of optimizers, as in the example below.
1optimizer = QuantizedSGD(model.parameters(), lr=0.01, momentum=0.9, exp_bits=5, sig_bits=10, rmode=1),
2
3optimizer = QuantizedRMSProp(model.parameters(), lr=0.01, exp_bits=5, sig_bits=10, rmode=5))
4
5optimizer = QuantizedAdagrad(model.parameters(), lr=0.01, exp_bits=5, sig_bits=10, rmode=4)
6
7optimizer = QuantizedAdam(model.parameters(), lr=0.001, exp_bits=5, sig_bits=10, rmode=6)
Customized Low precision optimization.
Pychop further provides the interface pychop.layers.post_quantization to convert model parameters into customized precision by specifying rounding and format parameters, enabling post-quantization simulation. We demonstrate the usage as below.
1from pychop.layers import post_quantization
2
3quantizer = LightChop(exp_bits=5, sig_bits=10, rmode=1) # define your customized fixed-point or floating point format
4quantized_model = post_quantization(model, quantizer)
Post quantization.
3.7.2. Straight-Through Estimator
The Straight-Through Estimator (STE) is a methodological framework widely used in training neural networks with discrete or non-differentiable operations, such as quantization or binarization. These operations challenge conventional backpropagation, which requires continuous gradients for parameter optimization. Non-differentiable functions, with their zero or undefined gradients, obstruct this process, impeding effective learning. The STE addresses this by approximating the gradient to enable training despite such discontinuities.
The STE operates by treating a non-differentiable function as differentiable during backward propagation. In the forward pass, it applies the intended discrete transformation, such as rounding a continuous value. In the backward pass, rather than using the operation’s true gradient—typically zero or undefined—it directly propagates the gradient from subsequent layers to preceding ones, bypassing the discrete step. This approximation allows gradient-based optimization to proceed, expanding the range of trainable neural architectures.
The Pychop framework integrates an STE module to support quantization seamlessly, enabling the neural network to adapt and learn despite the presence of discrete operations. Specifically, STE is leveraged to round activations to integer values during the forward pass while permitting gradients to propagate through during the backward pass as if the rounding operation had not occurred. This approach effectively reconciles the challenges posed by non-differentiable quantization, ensuring robust training of quantized neural networks.
3.8. Support for MATLAB
MATLAB provides built-in support for calling Python libraries through its Python Interface. This allows users to use Python functions, classes, and modules directly from MATLAB, making it easy to integrate Python-based scientific computing, machine learning, and deep learning libraries into MATLAB workflows. MATLAB interacts with Python by adding the py. prefix, which allows MATLAB to call the needed Python library seamlessly. One can also execute Python statements in the Python interpreter directly from MATLAB using the pyrun or pyrunfile functions. For detail, we refer the users to https://www.mathworks.com/help/matlab/call-python-libraries.html.
To trigger the Python virtual environment, one must have Python and the Pychop library installed (e.g., via pip manager using pip install pychop), then simply pass the following command in your MATLAB terminal:
1>> pe = pyenv() % or specify your python environment by ‘‘pe = pyenv(’Version’, ’/software/python/anaconda3/bin/python3’)‘‘
[…]
  PythonEnvironment with properties:
          Version: "3.10"
       Executable: "/software/python/anaconda3/bin/python3"
          Library: "/software/python/anaconda3/lib/libpython3.10.so"
             Home: "/software/python/anaconda3"
           Status: NotLoaded
    ExecutionMode: InProces
To use Pychop in your MATLAB environment, similarly, simply load the Pychop module:
1pc = py.importlib.import_module(’pychop’);
2ch = pc.LightChop(exp_bits=5, sig_bits=10, rmode=1)
3X = rand(100, 100);
4X_q = ch(X);
Or more specifically, use:
1np = py.importlib.import_module(’numpy’);
2pc = py.importlib.import_module(’pychop’);
3ch = pc.LightChop(exp_bits=5, sig_bits=10, rmode=1)
4X = np.random.randn(int32(100), int32(100));
5X_q = ch(X);
4. Simulations
4.1. Environmental settings
We run experiments on a Dell PowerEdge R750xa server [^5] https://front.convergence.lip6.fr/ with 2 TB of memory, Intel Xeon Gold 6330 processors (56 cores, 112 threads, 2.00 GHz), and an NVIDIA A100 GPU (80 GB HBM2, PCIe), providing robust computational power for large-scale simulations and deep learning tasks. We simulate the code in Python 3.10 and MATLAB R2024b. All simulations are performed on a single CPU and GPU.
Our analysis of image classification and object detection will primarily revolve around the following datasets:
  • MNIST(Deng, 2012): This dataset consists of handwritten digits ranging from 0 to 9, represented as grayscale images of size $28\times 28$ pixels. The dataset is widely regarded as a benchmark for image classification tasks and optical character recognition models. Each image is centered and normalized within a fixed-size frame, ensuring consistency across samples. The dataset contains 10 classes, corresponding to the 10 numerical digits, and exhibits variations in handwriting styles, stroke thickness, and digit positioning.
  • Fashion-MNIST(Xiao et al., 2017): The dataset consists of grayscale images, each of size $28\times 28$ pixels, depicting various fashion items such as clothing, footwear, and accessories. The dataset comprises 10 distinct classes, with each class representing a specific category of fashion products. The images are derived from the Zalando dataset and are designed as a more complex alternative to the MNIST dataset, introducing greater variability in shape and texture while preserving the same structural properties for benchmarking classification algorithms.
  • Caltech101(Li et al., 2022): The dataset contains $224\times 224$ RGB images of objects spanning 101 diverse categories, including animals, vehicles, and household items, with an additional background class. Collected by the California Institute of Technology, it presents a challenging benchmark for image classification due to its high intraclass variability and imbalanced sample sizes, ranging from 40 to 800 images per category, offering a robust testbed for evaluating generalization across heterogeneous visual patterns.
  • OxfordIIITPet(Parkhi et al., 2012): The dataset contains $224\times 224$ RGB images of pet animals across 37 breeds of cats and dogs, sourced from a collaboration between the University of Oxford and IIT. Designed to assess fine-grained classification, it features approximately 200 images per breed, split into training/validation and test sets, with significant variability in pose, lighting, and background, making it an ideal resource for studying detailed visual discrimination in real-world scenarios.
  • COCO(Lin et al., 2014): The COCO dataset comprises high-resolution color images of everyday scenes featuring objects from 80 distinct categories, including people, animals, vehicles, and household items. Originating from the Microsoft Common Objects in Context initiative, it offers a robust challenge beyond simpler datasets like ImageNet, with complex backgrounds, multiple objects per image, and annotations for both bounding boxes and segmentation masks. In our simulation code, we specifically utilize the COCO val2017 subset, which includes approximately 5,000 images from the validation split. This choice enables efficient evaluation of our quantized Faster R-CNN model’s performance on a diverse, well-annotated set without the computational overhead of the full training set (over 118,000 images) or the restricted access of the test set, ensuring rapid experimentation and reliable benchmarking of detection accuracy and latency trade-offs.
4.2. Speedup in MATLAB
Experimental simulations were conducted to compare the runtime performance of MATLAB’s chop function with Pychop for half-precision and bfloat16 precision rounding within the MATLAB environment. Additionally, Pychop’s performance was independently evaluated in a Python environment across various computational frameworks (NumPy and PyTorch) and hardware configurations (on CPU and GPU). The study assessed the baseline performance of MATLAB’s chop alongside Pychop, which implements the LightChop and Chop methods. Simulations tested square matrix sizes of $2{,}000$, $4{,}000$, $6{,}000$, $8{,}000$, and $10{,}000$, employing multiple rounding modes: round to nearest, round up, round down, round toward zero, and stochastic rounding. For clarity in the following discussion, MATLAB’s chop is denoted as mchop, while Pychop’s LightChop and Chop are referred to simply as LightChop and Chop, respectively.
The Figure 1 and 2. illustrate the runtime performance of the baseline mchop in comparison to Pychop, offering insights into scalability trends, framework efficiency, hardware influences, and optimization benefits. Results are presented in semilogarithmic plots, with distinct line styles distinguishing the data.
Although invoking Pychop within MATLAB introduces some runtime overhead, LightChop consistently outperforms mchop, while Chop on the CPU exhibits performance comparable to mchop. Furthermore, both Chop and LightChop achieve speedups of orders of magnitude over mchop when deployed on GPU hardware. Notably, the speedup ratio of LightChop becomes increasingly pronounced as the matrix size grows.
Figure 1. Runtime ratio of MATLAB ’s chop over Pychop in half precision (dashed for MATLAB-based, solid for Python-based).
Figure 2. Runtime ratio of MATLAB ’s chop over Pychop in bf16 precision (dashed for MATLAB-based, solid for Python-based).
4.3. Neural Network Quantization
Neural network quantization refers to applying reduced numerical precision (e.g., 16-bit floating-point, 8-bit integers, or even lower) in neural network training or inference instead of the standard 32-bit floating-point arithmetic typically used. In the following we explore the potential applications and advantages of utilizing Pychop in practical scenarios.
4.3.1. Image classifications
The image classification task in this study is simulated through a meticulously designed deep learning pipeline leveraging a pre-trained ResNet50 architecture (He et al., 2016), fine-tuned on datasets such as Caltech101 and OxfordIIITPet. The code employs a set of carefully selected hyperparameters to optimize model performance: a batch size of 64 balances computational efficiency and gradient stability, while a learning rate of 0.001, paired with the AdamW optimizer (weight decay = $1e^{-4}$), ensures robust convergence by adaptively adjusting parameter updates. Training spans 30 epochs, a duration sufficient to fine-tune the pre-trained weights— originally derived from ImageNet while mitigating overfitting, as evidenced by the cosine annealing learning rate scheduler (Loshchilov and Hutter, 2017) that gradually reduces the learning rate to refine optimization. Data augmentation techniques, including RandomCrop (Takahashi et al., 2018), RandomHorizontalFlip (Cubuk et al., 2020), RandAugment (Cubuk et al., 2020), and Cutout (DeVries and Taylor, 2017) (n_holes=1, length=32), enhance model generalization by introducing variability in the training samples, simulating real-world image distortions. The use of mixed precision training via PyTorch’s AMP (Automatic Mixed Precision) with a GradScaler accelerates computation and reduces memory demands without compromising accuracy. Furthermore, the incorporation of mixup data augmentation (alpha=1.0) (Zhang et al., 2018) and label smoothing (0.1) (Müller et al., 2019) in the Cross-entropy loss function further regularizes the model, encouraging robustness against noisy labels and overfitting. This configuration effectively simulates the image classification task by balancing feature extraction from pre-trained weights with task-specific adaptation, achieving high test accuracies over 90%, as validated through rigorous evaluation in both full-precision (fp32) training and inference phases of precision q43, q52, bfloat16, half, tf32, as well as three custom precisions with (5,5), (5,7), and (8,4) for exponent and significand bits. The results are depicted in Table 8, and the visualization of classification on Caltech101 are illustrated in Figure 3, Figure 4, Figure 5, and Figure 6, respectively.
In the analysis of accuracy across datasets, lower-precision float types like q43 and q52 generally underperform, achieving accuracies as low as 0.54% (q52, Caltech101, Round down) and rarely exceeding 12.26% (q43, MNIST, Round down), except for q52’s outlier of 99.50% on MNIST with round to nearest. In contrast, custom low-precision types—Custom 1 (5,5), Custom 2 (5,7), and Custom 3 (8,4)—consistently deliver high accuracies (e.g., 99.67% on MNIST, 92.63% on Caltech101), rivaling standard high-precision formats like half, bfloat16, tf32, and fp32, which stabilize at 91%–99.62% across all datasets and rounding methods. Notably, “Round to nearest” proves most reliable for maintaining accuracy across float types, yielding the highest average accuracies (e.g., 82.53% for FashionMNIST), while round toward zero mode occasionally boosts custom types. Thus, custom low-precision formats with 5–8 exponent bits and 4–7 significand bits can provide qualified accuracy (above 90%) for most tasks, offering an efficient trade-off between precision and performance.
Figure 3. Image classification on Caltech101 (Custom(5, 5)).
Figure 4. Image classification on Caltech101 (bfloat16).
Figure 5. Image classification on Caltech101 (tf32).
Figure 6. Image classification on Caltech101 (fp32).
Table 8. Accuracy Across Datasets and Float Types with Different Rounding Methods
Dataset Float Type Round to nearest Round up Round down Round toward zero Stochastic (prob.) Stochastic (uniform)
MNIST q43 9.17% 9.58% 12.26% 6.95% 7.58% 8.82%
q52 99.50% 9.58% 19.98% 99.16% 99.50% 99.27%
Custom 1 (5, 5) 99.62% 99.54% 99.66% 99.67% 99.64% 99.59%
Custom 2 (5, 7) 99.62% 99.60% 99.66% 99.64% 99.63% 99.62%
Custom 3 (8, 4) 99.63% 98.37% 99.47% 99.67% 99.65% 99.56%
half 99.62% 99.62% 99.62% 99.62% 99.62% 99.62%
bfloat16 99.62% 99.60% 99.66% 99.64% 99.63% 99.62%
tf32 99.62% 99.62% 99.62% 99.62% 99.62% 99.62%
fp32 99.62% 99.62% 99.62% 99.62% 99.62% 99.62%
Average 89.45% 90.57% 92.17% 89.29% 89.39% 89.48%
FashionMNIST q43 10.00% 10.00% 10.62% 10.00% 10.00% 10.00%
q52 91.41% 10.00% 29.51% 88.25% 90.32% 88.55%
Custom 1 (5, 5) 91.69% 90.03% 91.11% 91.85% 91.65% 91.78%
Custom 2 (5, 7) 91.60% 91.64% 91.83% 91.81% 91.63% 91.77%
Custom 3 (8, 4) 91.53% 85.90% 87.50% 91.69% 91.68% 91.60%
half 91.65% 91.65% 91.70% 91.68% 91.69% 91.66%
bfloat16 91.60% 91.64% 91.83% 91.81% 91.63% 91.77%
tf32 91.65% 91.65% 91.70% 91.68% 91.69% 91.66%
fp32 91.66% 91.66% 91.66% 91.66% 91.66% 91.66%
Average 82.53% 72.69% 75.27% 82.27% 82.44% 82.27%
Caltech101 q43 4.76% 7.14% 2.61% 2.38% 4.83% 4.76%
q52 89.41% 3.76% 0.54% 72.99% 84.27% 72.22%
Custom 1 (5, 5) 92.10% 82.81% 91.79% 92.56% 92.25% 92.40%
Custom 2 (5, 7) 92.56% 91.63% 92.33% 92.56% 92.63% 92.63%
Custom 3 (8, 4) 92.48% 56.41% 79.28% 92.17% 91.94% 91.63%
half 92.71% 92.40% 92.79% 92.71% 92.71% 92.63%
bfloat16 92.63% 91.63% 92.33% 92.56% 92.56% 92.63%
tf32 92.71% 92.40% 92.79% 92.71% 92.71% 92.63%
fp32 92.71% 92.71% 92.71% 92.71% 92.71% 92.71%
Average 82.45% 67.88% 70.80% 80.37% 81.85% 80.47%
OxfordIIITPet q43 2.73% 2.48% 2.70% 2.75% 2.75% 2.73%
q52 87.19% 2.73% 2.75% 74.79% 85.58% 62.85%
Custom 1 (5, 5) 91.14% 80.40% 87.54% 91.17% 90.98% 90.71%
Custom 2 (5, 7) 90.95% 90.57% 91.01% 91.01% 91.01% 91.01%
Custom 3 (8, 4) 90.84% 46.80% 70.48% 91.28% 90.71% 90.24%
half 91.09% 91.01% 90.95% 91.03% 91.09% 91.11%
bfloat16 90.95% 90.57% 91.01% 91.01% 91.01% 91.01%
tf32 91.09% 91.01% 90.95% 91.03% 91.09% 91.11%
fp32 91.09% 91.09% 91.09% 91.09% 91.09% 91.09%
Average 80.79% 65.18% 68.72% 79.46% 80.59% 77.98%
4.3.2. Object detection
Similar to our previous task, we employed a post-quantization approach to evaluate object detection performance on the COCO val2017 dataset, using various reduced-precision floating-point formats. Leveraging the Pychop library, we applied uniform low-precision conversion to all neural network parameters—through the LightChop class. This method enabled us to simulate a range of floating-point formats, incorporating six rounding modes: round to nearest, round up, round down, round toward zero, and stochastic rounding. To assess object detection accuracy, we used the mAP@0.5:0.95 metric, which averages the Average Precision (AP) across Intersection over Union (IoU) thresholds from 0.5 to 0.95. This metric evaluates both detection accuracy and localization precision by measuring how well predicted bounding boxes align with ground truth boxes at varying overlap levels, with results averaged across all classes and thresholds.
For this object detection task, which involves predicting object locations as bounding boxes defined by (x_min, y_min, width, height) alongside class labels and scores, we utilized Faster R-CNN (Ren et al., 2015) with a ResNet-50 Feature Pyramid Network (FPN) backbone (Lin et al., 2017). The model leverages pre-trained weights from the COCO dataset, accessible through PyTorch’s FasterRCNN_ResNet50_FPN_Weights.DEFAULT. This architecture integrates a ResNet-50 backbone for feature extraction, an FPN for multi-scale feature processing, a Region Proposal Network (RPN) for generating object proposals, and a detection head for bounding box regression and classification across 80 COCO categories plus a background class. The pre-trained weights stem from optimization on the COCO train2017 dataset ($\sim$118,000 images) over 12 epochs with a batch size of 2, employing a composite loss: the RPN combines binary cross-entropy loss for objectness classification with Smooth L1 loss for proposal regression, while the detection head uses cross-entropy loss for classification and Smooth L1 loss for box refinement, guided by a step-wise learning rate schedule (e.g., 0.02 initial rate, decayed at epochs 8 and 11). In our experiments, we applied these weights directly for inference on a subset of 100 images from COCO val2017, as specified by max_images=100, bypassing additional training. The post_quantization function from Pychop converted the model and its outputs into the specified precisions, allowing us to evaluate the impact of low-precision emulation on mAP@0.5:0.95 across the defined rounding strategies. The results are as depicted in Table 9 and visualized in Figure 7, Figure 8, and Figure 9, respectively.
In terms of the results, the choice of floating-point precision and rounding method significantly impacts performance, as measured by mAP@0.5:0.95. Low-precision formats like q43, q52, Custom (5,5), and Custom (5,7) consistently fail, delivering an mAP of 0.000 across all rounding methods. This suggests that their limited representational capacity—likely due to insufficient exponent or significand bits—renders them unusable for this task. In contrast, Custom (8,4) achieves a respectable mAP of 0.420 with “Round to nearest” peaking at 0.417 with stochastic (probabilistic) rounding. This marks it as the minimum precision capable of meaningful performance, far surpassing its low-precision counterparts. High-precision formats dominate the results. bf16 achieves a robust mAP of 0.423 with multiple rounding methods, while tf32 reaches a slightly higher peak of 0.424 with round up mode and fp32 maintains a steady 0.422 across all methods. These values, hovering around 0.42, indicate a precision threshold for object detection that exceeds typical requirements for image classification, where lower precision often suffices. Surprisingly, half precision fails entirely with an mAP of 0.000 across the board, hinting at potential implementation issues or an inability to handle the dynamic range and precision demands of bounding box predictions. Rounding methods also play a critical role. “Round to nearest” yields the highest average mAP@0.5:0.95 (0.210) across all formats, proving its reliability, while stochastic rounding methods (probabilistic at 0.209, uniform at 0.208) follow closely, offering competitive alternatives. Other methods—”Round up” (0.143), ”Round down” (0.162), and ”Round toward zero” (0.182)—lag behind, with ”Round up” performing the worst on average, likely due to systematic overestimation biases.
Table 9. mAP@0.5:0.95 by Floating Point Representation and Rounding Method
Round to nearest Round up Round down Round toward zero Stochastic (prob.) Stochastic (uniform)
q43 0.000 0.000 0.000 0.000 0.000 0.000
q52 0.000 0.000 0.000 0.000 0.000 0.000
Custom (5, 5) 0.000 0.000 0.000 0.000 0.000 0.000
Custom (5, 7) 0.000 0.000 0.000 0.000 0.000 0.000
Custom (8, 4) 0.420 0.026 0.197 0.377 0.417 0.415
bfloat16 0.423 0.419 0.423 0.416 0.423 0.420
tf32 0.422 0.424 0.420 0.422 0.421 0.422
half 0.000 0.000 0.000 0.000 0.000 0.000
fp32 0.422 0.422 0.422 0.422 0.422 0.422
Average 0.210 0.143 0.162 0.182 0.209 0.208
Figure 7. Object detection using bfloat16 (Red: Ground Truth, Green: Predictions).
Figure 8. Object detection tf32 (Red: Ground Truth, Green: Predictions).
Figure 9. Object detection using fp32 (Red: Ground Truth, Green: Predictions).
5. Conclusion and future work
In this work, we present Pychop, an open-source software designed as an efficient precision emulation tool for numerical methods and deep learning research. By integrating seamlessly with automatic differentiation frameworks such as PyTorch, JAX, and NumPy, Pychop bridges theoretical exploration and practical deployment, enhancing accessibility and usability in computational science. Its flexibility, comprehensive rounding support, and rigorous validation establish it as a vital resource for advancing mixed-precision numerical algorithms and deep learning applications.
To evaluate its adaptability and real-world utility, we employed Pychop to simulate post-quantization effects in image classification and object detection tasks across established datasets. These experiments provide insights into the optimal bitwidths for exponents and significands required for high-quality inference, revealing performance trade-offs to inform the selection of efficient floating-point representations for specific use cases.
Pychop is poised for continuous development and innovation. While currently focused on neural network quantization and mixed-precision computing, we plan to extend compatibility to additional frameworks, such as TensorFlow, to amplify its impact and adoption. We invite contributions from the research and development community—via GitHub pull requests and issue submissions—to enhance its functionality, optimize performance, and explore novel applications, ensuring Pychop remains a robust and versatile tool for future advancements.